HUBBLE SPACE TELESCOPE DAILY REPORT #2533 PERIOD COVERED: 0000Z (UTC) 01/05/00 - 0000Z (UTC) 01/06/00 Daily Status Report as of 006/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of STIS/CCD 8530 (SMOV3a STIS BEA Science) The Space Telescope Imaging Spectrograph (CCD) was used to obtain SMOV3a Bright Earth Avoidance time period for science exposures with the STIS CCD. Various kinds of targets are observed. The proposal completed nominally. 1.2 Completed WF/PC-2 8497 (Point Spread Function Verification) The WF/PC-2 was used to verify the WFPC2 PSF after SM3A. There were no reported problems. 1.3 Completed Three Sets of FGS 8532 (White Dwarf and M-Stars in the Solar Neighborhood; Revealing Binarity) FGS-1 was used to conduct a high angular resolution survey with FGS1r to detect binarity among nearby M dwarfs and white dwarfs. The first iteration of this proposal was affected by the failure to perform astrometry as reported in 2.1 and HSTAR 7463. Otherwise, the observations completed with no further reported problems. 1.4 Completed STIS/CCD 8507 (CCD Darks) The Space Telescope Imaging Spectrograph (CCD) was used to obtain weekly updates to the hot pixel list. The proposal completed as planned. 1.5 Completed STIS/CCD 8508 (CCD Bias Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to obtain updates to the weekly superbiases. The proposal was completed as planned. 1.6 Completed STIS/MA2 8509 (NUV MAMA Darks) The Space Telescope Imaging Spectrograph (MA2) was used to determine the health of the FUV MAMA detector by monitoring the background count rate. The observations completed with no reported problems. 1.7 Completed WF/PC-2 8492 (UV Monitor for -88C Cooldown) The WF/PC-2 was used to conduct the UV monitor after the first cooldown to -88C. Throughput measurements are also made using UV filters and UV crossed with long-pass filters to determine red leak contribution. -These are used to indicate any drop in throughput at Lyman alpha, which could be due to contamination in the pick-off mirror. The proposal completed nominally. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 15 Successful: 15 As detailed in HSTAR 7463, while the acquisition at 005/185056Z succeeded, the astrometry failed. The proposal described in 1.3 was affected. HSTAR 7465 similarly describes failed astrometry occurring at 006/001200Z. An FGS proposal , appearing during the period of the next report, was affected. Scheduled Re-acquisitions: 0 Successful: 0 2.2 FHST Updates: Scheduled: 43 Successful: 43 2.3 Operations Notes: The Servicing Mission Orbital Verification (SMOV) process continues. ROP SR-1A was utilized to clear the SSR EDAC errors three times. "Emergency" CCL procedure updates were performed on all operational CCS strings at 005/1445Z. An operations request was performed to obtain the location of the multi-bit NICMOS error at 005/1505Z. Subsequently, NICMOS was successfully recovered from "safe" to "boot" at 005/1524Z. 486 software patch 1.2 was successfully installed at 005/2157Z. This patch modifies the Background Memory Integrity Check, alters the internal tables to allow the SSR internal relay timer to be changed, changes code to correct its handling of a four-battery system, changes the size of the FGE refresh tables, and modifies the Battery Pressure Safemode Limit Changes. HSTAR 7464 documents a STIS EMC retry at 005/205750Z. ROP NS-12 was executed at 005/2221Z to clear the STIS flight software error flags. A 486 table was dumped at 005/2305Z per ROP DF-1A. An SSR-1 retry/interrupt log patch load was installed at 006/0258Z with no reported problems. The patch was subsequently tested successfully. Using ROP IC-2, the SSA transmitter was turned on and off at 006/0505Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of SMOV and normal science observations and calibrations. /CAW